OpenStudy (anonymous):

need some help with a calculus question dealing with Integration by parts. ∫ln(3x-2)dx

OpenStudy (amistre64):

(3x-2) ln(3x-2) - (3x-2) i wonder

OpenStudy (anonymous):

Integrate by parts will work on this..

OpenStudy (anonymous):

what would be u and what would be dv though?

OpenStudy (anonymous):

\[\int\ln(x)dx=x\ln(x)-x\] by parts, memorize it and adjust as needed

OpenStudy (amistre64):

ln(x) ints up into: xlnx - x + c

OpenStudy (anonymous):